The approximate number of tourists in South America is 40.3million and its average rate of change is an increase of 0.8million per year.Also the approximate number of tourists in the middle East is 17.0million and its average rate of change is an increase of 1.8million per year. A-Define the variables and write an equation for each regions tourism rate.
B-If the trends continue, in how many years would you expect the number of tourists in the regions be equal

Answer by Rongtian(7)

A)let x=years. A=40.3+0.8x E=17+1.8x
40.3+0.8x=17+1.8x
23.3=x
23.3 years